# Hypergraph Representations of scRNA-seq Data for Improved Clustering with Random Walks

### v3

@@ Abstract
Analysis of single-cell RNA sequencing data is often conducted through considering their network projections such as co-expression networks, primarily due to the abundant availability of network analysis tools for downstream tasks. However, this approach has several limitations. First, projecting scRNA-seq data into networks can result in the loss of higher-order information since it summarizes the gene expression profiles across cells into pairwise cell or gene similarities. Secondly, scRNA-seq data is often very sparse, but the constructed co-expression network projections are fully connected without sparsification, resulting in dense and inefficient data representation. Further, the sparsity of scRNA-seq data leads to co-occurrence of zero expression in a large proportion of the genes. Such zero-expression co-occurence will be accounted as similarity in gene expression among cells by correlation methods such as Spearman’s rank and Pearson’s correlation coefficient and lead to an overestimation of co-expression.
To address these limitations, we propose conceptualizing scRNA-seq expression data as hypergraphs, which are generalized graphs in which the hyperedges can connect more than two vertices. This flexibility makes hypergraphs particularly well-suited for representing scRNA-seq data without requiring further transformation such as co-expression network generation. In the context of scRNA-seq data, the hypergraph nodes represent cells and the edges represent genes. Specifically, each hyperedge connects to all the cells where the gene is actively expressed and records the expression of the gene across different cells. The hypergraph conceptualization enables us to explore multi-way relationships beyond the pairwise interactions in co-expression networks for scRNA-seq data analysis while utilizing network analysis tools.
We propose two novel hypergraph random walk-based cell clustering methods: (1) the Dual-Importance Preference Hypergraph Walk (DIPHW) and (2) the co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW). These methods outperform widely used and established methods in terms of clustering accuracy on both simulated and real scRNA-seq datasets. The improvement is especially significant when data modularity is weak. By conceptualizing the scRNA-seq data as hypergraphs, our methods provide a more nuanced representation of gene expression data, allowing the incorporation of network perspectives without the information loss caused by enforcing a unipartite co-expression network projection. Furthermore, CoMem-DIPHW incorporates the gene co-expression network, cell co-expression network, and the cell-gene expression hypergraph from the single-cell abundance counts data altogether for embedding computation. This approach accounts for both the local information from single-cell level gene expression and the global information from pairwise similarities in the two co-expression networks.
@@ I Introduction
Network analysis has become a popular tool for studying complex systems in biology, as a reflection of the combinatorial interaction mechanism of biomolecules. Its applications include functional analysis  [1, 2, 3, 4] and interaction prediction in protein-protein interaction (PPI) networks [5, 6, 7], regulator identification and pathway analysis in gene regulatory networks [8, 9, 10], disease modeling, prediction, and intervention in epidemiology [11, 12, 13], and cell type identification in co-expression networks [14, 15, 16, 17].
However, traditional methods represent the biology systems using unipartite networks, which only captures relationships (edges) among one type of entity (nodes), are often not the most natural or information-preserving representation of the complex systems since many involve interactions between multiple types of entities. Instead of modelling these systems using network models that could capture higher order interactions, most often, a unipartite network projection is chosen such that the system could be analyzed with the abundant network analysis methods developed based on unipartite networks [23, 14]. For example, in protein-protein interaction (PPI) networks, unipartite representations capture relationships between proteins but fail to capture the biological functions associated with groups of proteins. Similarly, unipartite projections of gene regulatory networks connect genes based on regulatory relationships but do not capture relationships among different entities such as transcription factors or their binding sites.
Single-cell RNA sequencing (scRNA-seq) has enabled the profiling of gene expression at the individual cell level [24, 25, 26, 27], whereas conventional bulk tissue RNA sequencing measures expression at the tissue level, averaging gene expression over an ensemble of cells. Cell type identification, as one of the most important downstream tasks in single-cell RNA-seq (scRNA-seq) data analysis  [28], has applications across biology and medicine, including tracing the trajectories of distinct cell lineages in development for cell differentiation studies  [29], tissue heterogeneity analysis for cancer research  [30, 31, 32], immune cell profiling for therapy development  [33, 34, 35, 36], and biomarker discovery for diagnosis.
@@
Given the issues with unipartite projections, it is crucial to explore alternative network representations that can accurately reflect the scRNA-seq data. Hypergraphs offer a neat solution by representing multiway relationships in scRNA-seq directly, without the need for further data projection.
Figure 1: Network Representations of the scRNA-seq Data. The scRNA-seq cell-gene expression data can be represented by various networks, including unipartite network projections, such as cell and gene co-expression networks. In a cell co-expression network, the nodes are cells and the edge weights represent similarity in cell-pair expression. Similarly, in gene co-expression networks, nodes represent genes. Additionally, hypergraph and bipartite network representations preserve the exact expression level of a gene in a cell. In the hypergraph representation, nodes represent cells and hyperedges represent genes. In the bipartite representation, one set of nodes represents cells, and the other set represents genes.
Figure 2: CoMem Random Walk Illustration. In the hypergraph representation of the scRNA-seq data, nodes represent cells and edges represent genes. The expression level of a gene in a cell is captured by the weight of a node in an edge. Random walks are performed on the hypergraph by alternately considering the probability of the walker choosing a node given an edge, and choosing an edge given the current node. For DIPHW, the input is the hypergraph representation of the scRNA-seq data. For CoMem-DIPHW, the input includes the hypergraph, cell co-expression network, and gene co-expression network, which serve as node and edge similarity matrices to aid memory incorporation during the walk. A small neural network (Word2Vec) is applied to the sampled node paths from the hypergraph random walk process to compute cell embeddings. Finally, cell clustering is achieved by applying K-means to the obtained cell embeddings.
@@ II Methods
We conceptualize scRNA-seq data as a hypergraph $\mathcal{H}$ and design two hypergraph random walk algorithms that capture cell-cell relationships by sampling the cell-to-gene and gene-to-cell interactions in the conceptualized hypergraph. We emphasize that this hypergraph representation of the scRNA-seq data does not require further transformation of the data and consequently would not result in any information loss or additional computational cost.
@@ II.1 Hypergraph representation
The single-cell RNA sequencing (scRNA-seq) data can be conceptualized as a hypergraph $\mathcal{H}=(V,E)$, where:
  • $V$ represents the vertex set corresponding to the cells profiled in the experiment.
  • $E$ denotes the hyperedge set representing genes, with each hyperedge $e\in E$ connecting to all cells $\in V$ in which the gene $e$ is actively expressed. The weight of connection of a hyperedge (gene) to each node (cell) is determined by the gene’s expression level in that cell.
@@
$V$ Set of all nodes (cells)
$E$ Set of all hyperedges (genes) in the hypergraph
$G_{V}$ Cell co-expression network/ Node similarity matrix
$G_{E}$ Gene co-expression network/ Edge similarity matrix
$\Gamma_{e}(v)$ Weight of node $v$ in hyperedge $e$, i.e., expression of gene $e$ in cell $v$
@@
$P(V_{t+1}=w\mid E_{t+1}=e,V_{t}=v)$ Probability of selecting node $w$ at time $t+1$, given hyperedge $e$ and previous node $v$
@@
Table 1: Hypergraph Random Walk Notation Table.
@@ II.2 Method 1: Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW)
Building on this hypergraph representation, we introduced a novel hypergraph random walk approach for cell embedding computation, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW), which accounts for both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering. In a hypergraph random walk process, the walker alternates between transitioning from a hyperedge to a node and from a node to a hyperedge, following the node-to-edge transition probabilities $P_{E|V}(e,v)$ and edge-to-node transition probabilities $P_{V|E}(v,e)$. In the hypergraph random walk with edge-dependent vertex weight (EDVW) by Chitra and Raphael (2019) [41], the edge-to-node transition probability is edge-dependent, but the node-to-edge transition probability does not consider the importance of the edge to the node. As a natural extension to EDVW, we defined the node-to-edge transition probability to be vertex-dependent in DIPHW, and used a preference exponent to accelerate the clustering process. The modified random walk process is described as follows:
@@
    Node-to-Hyperedge Transition Probability:
@@
    Hyperedge-to-Node Transition Probability:
@@
    Node-to-Node Transition Probability:
    The unipartite node-to-node transition probability of DIPHW is obtained by summing the product of the node-to-edge and edge-to-node transition probabilities over all hyperedges. This probability is used to sample random walk node paths, which serve as inputs to a neural network for cell embedding computation.
@@
    Node-to-Node Transition Probability in Vectorized Format:
    The $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ could be written in the vectorized form as,
@@
    Vectorization eliminates explicit iteration and reduces the computation time. Here, $W_{E|V}=\mathbf{I}_{\mathcal{H}}^{T}W_{E}$ is the $|V|\times|E|$ node-to-edge transition weight matrix where column $e$ of $W_{E|V}$ contains the vertex weights in hyperedge $e$ multiplied by the weight of hyperedge $e$. Similarly, $W_{V|E}$ represents the edge-to-node transition weight matrix before normalization with dimensions $|E|\times|V|$, given by $W_{V|E}=\mathbf{I}_{\mathcal{H}}^{\alpha},\quad w_{ve}=\left[\mathbf{I}_{\mathcal{H}}(e,v)\right]^{\alpha}$. Finally, $D_{E|V}$ and $D_{V|E}$ are diagonal matrices that normalize vertex-to-edge and edge-to-vertex transitions to ensure the probability distribution sums to 1.
@@
We now introduce CoMem-DIPHW, an extension of DIPHW that incorporates co-expression networks and single-cell transcriptomic profiles. The memory mechanism of CoMem-DIPHW leverages cell and gene co-expression networks to account for previously visited nodes and edges during transitions to capture both local expression relationships and global co-expression patterns.
@@
We remark that a matrix can be directly interpreted as the incidence matrix of a hypergraph to represent the relationship between two variables, without additional computation.
@@
Using scRNA-seq data, we construct two co-expression networks: the cell-co-expression network (CCN) $G_{V}$ and the gene-co-expression network (GCN) $G_{E}$. In the CCN, cells are nodes, and edges correspond to correlation coefficients between cell pairs, indicating the similarity between their expression profiles. Similarly, in the GCN, genes are nodes and edges correspond to correlation coefficients between gene pairs. Spearman’s rank coefficient was chosen to measure pairwise similarity of expression profiles as it captures non-linearity better than Pearson’s correlation. These unipartite networks serve as node and edge similarity matrices, embedding global-level information that summarizes associations between cell and gene pairs.
@@
CoMem-DIPHW integrates a memory component by considering the similarity between consecutively visited nodes and edges. This memory mechanism adjusts the transition probabilities such that the node-to-edge transition probability depends on the previous edge visited, and the edge-to-node transition probability depends on the previous node visited. The node-node and edge-edge dependencies are incorporated through the node and edge similarity networks, which are the two correlation networks derived from the transcriptomic expression data. The memory mechanism constrains the hypergraph walker to focus on cell-gene interactions among similar nodes(cells) and edges(genes) to better uncover modularity within the data.
While the co-expression networks provide global perspectives through unipartite projections, specific expression levels of genes in individual cells (local information) are preserved in the hypergraph structure. CoMem-DIPHW integrates both global and local information for more accurate cell type identification.
@@
The random walk process in CoMem-DIPHW is formulated as follows:
Let $G_{V}$ be the vertex similarity matrix, where $G_{V}(v,w)$ quantifies the similarity between vertices $v$ and $w$. Similarly, let $G_{E}$ be the edge similarity matrix, where each element $G_{E}(e,f)$ represents the similarity between edge $e$ and $f$.
@@
To encourage exploration and ensure non-lazy walks, we modify the node and edge similarity graphs such that the diagonal elements are set to zero. This ensures that during the random walk, the random walker does not remain at the same edge or node in consecutive steps.

### v4

# Hypergraph Representations of Single-Cell RNA Sequencing Data for Improved Cell Clustering

@@ Abstract
Abstract
Motivation:
Single-cell RNA sequencing (scRNA-seq) data analysis is often performed using network projections that produce co-expression networks. These network-based algorithms are attractive because regulatory interactions are fundamentally network-based and there are many tools available for downstream analysis. However, most network-based approaches have two major limitations. First, they are typically unipartite and therefore fail to capture higher-order information. Second, scRNA-seq data are often sparse, so most algorithms for constructing unipartite network projections are inefficient and may overestimate co-expression relationships, or may under-utilize the sparsity when clustering (e.g., with cosine distance). To address these limitations, we propose representing scRNA-seq expression data as hypergraphs, which are generalized graphs where a hyperedge can connect more than two nodes. In this context, hypergraph nodes represent cells, and hyperedges represent genes. Each hyperedge connects all cells in which its corresponding gene is actively expressed, indicating the expression of that gene across different cells. The resulting hypergraph can capture higher-order information and appropriately handle varying levels of data sparsity. This representation enables clustering algorithms to leverage higher-order relationships for improved cell-type differentiation.
Results: To distinguish cell types using hypergraph representations of scRNA-seq data, we introduce two novel clustering algorithms: (1) Dual-Importance Preference Hypergraph Walk (DIPHW) and (2) Co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW). DIPHW is a new hypergraph-based random walk algorithm that computes cell embeddings by considering the relative importance of genes to cells and cells to genes, incorporating a preference exponent to facilitate clustering. CoMem-DIPHW integrates two unipartite projections, the gene co-expression and cell co-expression networks, along with the cell-gene expression hypergraph derived from single-cell abundance count data into the random walk model. The advantage of CoMem-DIPHW is that it accounts for both local information from single-cell gene expression and global information from pairwise similarity in the two co-expression networks. We benchmark the performance of our algorithms against established and state-of-the-art deep learning approaches using both real-world and simulated scRNA-seq data. Real-world datasets include cells from the human pancreas, mouse pancreas, human brain, and mouse brain tissues. We also use a ground-truth labeled cell-type annotation dataset based on human lung adenocarcinoma cell lines. Quantitative evaluation shows that CoMem-DIPHW consistently outperforms established algorithms and state-of-the-art deep learning algorithms for cell-type clustering. Our proposed algorithms show the greatest improvement on scRNA-seq data with weak modularity. Moreover, CoMem-DIPHW successfully annotates clusters with biologically relevant cell types. Our results highlight the utility of hypergraph representations in the analysis of scRNA-seq data.
Availability and Implementation: Our methods are implemented in Python and are available at https://github.com/wanhe13/CoMem-DIPHW.
@@ I Introduction
Network analysis has become a popular tool for studying complex systems in biology, reflecting the combinatorial interactions of biomolecules. Its applications include functional analysis [szklarczyk2015string, veres2015comppi, vazquez2003global, ashtiani2018systematic] and prediction of interactions in protein-protein interaction (PPI) networks [kovacs2019network, lei2013novel, murakami2017network], identification of regulators and analysis of pathways in gene regulatory networks [coleman2023gene, muzio2021biological, zickenrott2016prediction], disease modeling, prediction, and intervention in epidemiology [kraemer2020effect, chinazzi2020effect, davis2021cryptic], and identification of cell types in co-expression networks [langfelder2008wgcna, satija2015spatial, butler2018integrating, wolf2018scanpy].
However, traditional methods represent biological systems using unipartite networks, which capture relationships only between a single type of entity. Such representations are often not the most natural or information-preserving choice, since many complex systems involve interactions between multiple types of entities. Most often, a unipartite network projection is chosen such that the system could be analyzed with the abundant network analysis methods developed based on unipartite networks [horvat2012one, langfelder2008wgcna]. However, this focus on unipartite networks restricts our attention to biological interactions just between pairs of genes, proteins, or other kinds of nodes. [^1] We use the terms node and vertex interchangeably. In reality, biological processes often entail multiway interactions that cannot be represented by unipartite networks. For instance, three-gene epistatic interactions vastly outnumber pairwise epistasis in experimental analyses of yeast genetic networks [kuzmin2018systematic]. In protein-protein interaction (PPI) networks, unipartite representations capture relationships between proteins but fail to capture the biological functions associated with groups of proteins. Similarly, unipartite projections of gene regulatory networks connect genes based on regulatory relationships but do not capture relationships among different entities such as transcription factors or their binding sites. Therefore, network analysis methods are needed to more accurately capture the higher-order interactions that are apparently ubiquitous in biological systems.
Single-cell RNA sequencing (scRNA-seq) has enabled the profiling of gene expression at the individual cell level [yanai2019cel, macosko2015highly, zheng2017massively, picelli2014full], whereas conventional bulk tissue RNA sequencing measures expression at the tissue level, averaging gene expression over an ensemble of cells. Cell type identification, as one of the most important downstream tasks in single-cell RNA-seq (scRNA-seq) data analysis [hwang2018single], has applications in biology and medicine, including tracing the trajectories of different cell lineages in the development of cell differentiation studies [marioni2017single], tissue heterogeneity analysis for cancer research [wong2022single, laviron2022tumor, karaayvaz2018unravelling], immune cell profiling for therapy development [landhuis2018single, lyons2017immune, ermann2015immune, chung2017single], and biomarker discovery for diagnosis.
@@
Given the issues with unipartite projections, it is crucial to explore alternative network representations that can accurately reflect the scRNA-seq data and the underlying biological reality of higher-order interactions among genes. Hypergraphs [bianconi2021higher, battiston2020networks, battiston2021physics] offer a nice solution by directly representing multiway relationships in scRNA-seq data, without requiring further data projection. Existing work on hypergraph representations has demonstrated the improved performance they can achieve in modeling complex biological systems by incorporating higher-order interactions. For example, Liu et al. [liu2022multi] developed a hypergraph-based neural network to predict synergistic drug combinations for cancer treatment. Wang et al. [wang2024hypertmo] integrated multi-omics data with hypergraph convolutional neural networks to classify patients with diseases such as breast cancer and Alzheimer’s. Gaudelet et al. [gaudelet2018higher] used hypergraphs to model interactions across different levels of protein organization (such as protein-protein interactions, complexes, and pathways) to better capture the complexity of biological systems and predict biological functions. Ma and Ma [ma2022hypergraph] developed a hypergraph-based logistic matrix factorization method that predicts possible metabolite–disease interactions and uncovers novel disease-related metabolites. These successes highlight the potential of hypergraphs to model multiway relationships inherent in biological systems, motivating their application to scRNA-seq analysis.
Figure 1: Network Representations of scRNA-seq Data. Expression data can be represented by various networks, including unipartite network projections such as cell and gene co-expression networks. In a cell co-expression network, nodes are cells and edge weights represent similarity in cell-pair expression. In gene co-expression networks, nodes represent genes. Hypergraph and bipartite network representations preserve the exact expression level of a gene in a cell. In the hypergraph representation, nodes represent cells and hyperedges represent genes. In the bipartite representation, one set of nodes represents cells and the other set represents genes.
Figure 2: Hypergraph Clustering Algorithm Illustration. In the hypergraph representation of scRNA-seq data, nodes represent cells and edges represent genes. The expression level of a gene in a cell is captured by the weight of a node in an edge. Random walks are performed on the hypergraph by alternately considering the probability of the walker choosing a node given an edge and choosing an edge given the current node. For DIPHW, the input is the hypergraph representation of the scRNA-seq data. For CoMem-DIPHW, the input includes the hypergraph, cell co-expression network, and gene co-expression network, which serve as node and edge similarity matrices to incorporate memory during the walk. A small neural network (word2vec) is applied to the sampled node paths from the hypergraph random walk process to compute cell embeddings. Finally, cell clustering is performed by applying K-means to the resulting cell embeddings.
@@ II Methods
We conceptualize scRNA-seq data as a hypergraph $\mathcal{H}$ and design two hypergraph random walk algorithms that capture cell-cell relationships by sampling cell-to-gene and gene-to-cell interactions in the hypergraph. This hypergraph representation of scRNA-seq data does not require additional data transformation, does not cause information loss, or increase computational cost.
@@ II.1 Hypergraph representation
scRNA-seq data can be conceptualized as a hypergraph $\mathcal{H}=(V,E)$, where:
  • $V$ represents the set of nodes corresponding to the cells profiled in the experiment.
  • $E$ denotes the set of hyperedges representing genes, with each hyperedge $e\in E$ connecting to all cells $\in V$ in which the gene $e$ is actively expressed. The weight of connection of a hyperedge (gene) to each node (cell) is determined by the gene’s expression level in that cell.
@@
$V$ Set of nodes (cells)
$E$ Set of hyperedges (genes) in the hypergraph
$G_{V}$ Cell co-expression network (node similarity matrix)
$G_{E}$ Gene co-expression network (edge similarity matrix)
$\Gamma_{e}(v)$ Weight of node $v$ in hyperedge $e$ (i.e., expression of gene $e$ in cell $v$)
@@
$P(V_{t+1}=w\mid E_{t+1}=e,V_{t}=v)$ Probability of selecting node $w$ at time $t+1$, given hyperedge $e$ at time $t+1$ and previous node $v$
@@
Table 1: Notation Table for Our Hypergraph Random Walk Representation.
@@ II.2 Method 1: Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW)
Building on this hypergraph representation, we introduce a novel hypergraph random walk approach for cell embedding computation, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW), which accounts for both the relative importance of edges to nodes and nodes to edges, or the importance of genes relative to cells and cells to genes, along with a preference exponent for faster clustering. In a hypergraph random walk process, the walker alternates between transitioning from a hyperedge to a node and from a node to a hyperedge, following the node-to-edge transition probabilities $P_{E|V}(e,v)$ and edge-to-node transition probabilities $P_{V|E}(v,e)$. In the hypergraph random walk with edge-dependent vertex weight (EDVW) by Chitra and Raphael (2019) [Chitra2019], the edge-to-node transition probability is edge-dependent, but the node-to-edge transition probability does not consider the importance of the edge to the node. As a natural extension to EDVW, we define the node-to-edge transition probability to be vertex-dependent in DIPHW and use a preference exponent to accelerate the clustering process. The modified random walk process is as follows:
@@
    Node-to-Hyperedge Transition Probability
@@
    Hyperedge-to-Node Transition Probability
@@
    Node-to-Node Transition Probability
    The unipartite node-to-node transition probability of DIPHW is calculated by summing the products of the node-to-edge and edge-to-node transition probabilities across all hyperedges. This probability is used to sample random walk node paths, which are then input into a neural network to compute cell embeddings.
@@
    Node-to-Node Transition Probability in Vectorized Format
    The $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ could be written in the matrix form as
@@
    The matrix formulation eliminates explicit iteration and reduces computation time. Here, $W_{E|V}=\mathbf{I}_{\mathcal{H}}^{T}W_{E}$ is the $|V|\times|E|$ node-to-edge transition weight matrix, where column $e$ of $W_{E|V}$ contains the vertex weights in hyperedge $e$ multiplied by the weight of hyperedge $e$. Similarly, $W_{V|E}$ represents the edge-to-node transition weight matrix before normalization with dimensions $|E|\times|V|$, given by $W_{V|E}=\mathbf{I}_{\mathcal{H}}^{\alpha},\quad w_{ve}=\left[\mathbf{I}_{\mathcal{H}}(e,v)\right]^{\alpha}$. Finally, $D_{E|V}$ and $D_{V|E}$ are diagonal matrices that normalize vertex-to-edge and edge-to-vertex transitions to ensure the probability distribution sums to 1.
@@
CoMem-DIPHW is an extension of DIPHW. It incorporates co-expression networks and single-cell transcriptomic profiles. The memory mechanism of CoMem-DIPHW leverages cell and gene co-expression networks to account for previously visited nodes and edges during transitions to capture both local expression relationships and global co-expression patterns.
  1. 1.
@@
    We note that a matrix can be directly interpreted as the incidence matrix of a hypergraph to represent the relationship between two variables, without additional computation.
  2. 2.
@@
    Using scRNA-seq data, we construct two co-expression networks: the cell co-expression network ($G_{V}$) and the gene co-expression network ($G_{E}$). In the cell co-expression network, cells are nodes, and edges represent correlation coefficients between cell pairs, indicating the similarity of their expression profiles. Similarly, in the gene co-expression network, genes are nodes, and edges represent correlation coefficients between gene pairs. We use Spearman’s rank coefficient to measure pairwise similarity of expression profiles, as it captures non-linearity better than Pearson’s correlation. These unipartite networks serve as node and edge similarity matrices, embedding global-level information that summarizes associations between cell and gene pairs.
  3. 3.
@@
    CoMem-DIPHW incorporates a memory component by considering the similarity between consecutively visited nodes and edges. This memory mechanism adjusts the transition probabilities so that the node-to-edge transition probability depends on the previously visited edge, and the edge-to-node transition probability depends on the previously visited node. Node-node and edge-edge dependencies are incorporated through the node and edge similarity networks, which are two correlation networks derived from transcriptomic expression data. The memory mechanism constrains the hypergraph walker to focus on cell-gene interactions among similar nodes (cells) and edges (genes), thus better revealing modularity within the data.
    While the co-expression networks provide global perspectives through unipartite projections, the specific expression levels of genes in individual cells (local information) are preserved in the hypergraph structure. CoMem-DIPHW integrates global and local information for more accurate cell type identification.
  4. 4.
@@
The random walk process in CoMem-DIPHW is formulated as follows. Let $G_{V}$ be the vertex similarity matrix, where $G_{V}(v,w)$ quantifies the similarity between nodes $v$ and $w$. Similarly, let $G_{E}$ be the edge similarity matrix, where each element $G_{E}(e,f)$ represents the similarity between edge $e$ and $f$.
@@
In CoMem, we calculate the unipartite node-to-node transition probability of the hypergraph random walker. To do this, we first calculate the following four probabilities.
@@
To encourage exploration and ensure non-lazy walks, we modify the node and edge similarity graphs such that the diagonal elements are set to zero. This ensures that during the random walk, the random walker does not remain at the same node or edge in consecutive steps.
@@
II.4 Cell Embedding and Clustering
We perform random walks using the defined hypergraph node-to-edge and edge-to-node transition probabilities to sample sequences of visited cells. These sampled cell paths, which capture the geometry of the hypergraph representing the scRNA-seq data, are then used as input to word2vec [mikolov2013distributed] to learn low-dimensional vector representations of cells. Cells with similar expression profiles frequently co-occur in the sampled random walk paths and are closer in the embedded space. Finally, K-means clustering is applied to cell embeddings to obtain cluster assignments.
II.5 scRNA-seq Data Simulation
In our study, we develop an algorithm to simulate data that mimics the characteristics of scRNA-seq expression data (see Algorithm 1). In addition to the underlying assumption of within-cell-type homogeneous expression that governs most cell clustering methods, we incorporate between-cell-type crosstalk [armingol2021deciphering, zhang2023defining, zepp2019cellular, bayik2021cancer, lehuen2010immune] by integrating configurable intermodular covariance, density, and signal strength into the simulation model. The algorithm generates a sparse matrix representing the simulated scRNA-seq expression data, with numerous user-configurable parameters, including number of modules, module density, shape, background signal strength, modular signal strength, intermodular signal strength, intermodular covariance, and noise level (see Table S4). This method is implemented as a function in our codebase and is available for use by other researchers [wan2025comem].
Our simulation model is highly flexible to accommodate various types of scRNA-seq data. For example, the model can simulate both sequenced datasets with high density (up to 50%) and sequenced datasets with low density (as low as 1%) [andrews2021tutorial]. The proportion of non-housekeeping genes can be adjusted by changing the number of differentially expressed genes (DEGs) in each module. In addition, the variance in module sizes and the number of co-expressed genes in each cell type can be configured by setting different variances for the Poisson distribution used in the embedded module simulation. The modules embedded in the simulated data can be configured with higher density and average signal strength to model the expression profiles of biomarker genes that are highly expressed in certain cell types. This enables us to mimic the sparsity and modularity of real scRNA-seq data, where a small number of genes are highly expressed in specific cell types. Alternatively, the embedding modules can be user-specified without using the embedded module simulation.
Algorithm 1 Simulation Algorithm for scRNA-seq Data
1:Input: Number of genes $E$, number of cells $V$, number of cell types $K$; average genes per cell type $\bar{g}$ and average cells per cell type $\bar{c}$; average expression levels for within-cell-type $\lambda_{\text{ct}}$, cross-cell-type $\lambda_{\text{cross}}$, and background $\lambda_{\text{bg}}$; densities within-cell-type $\rho_{\text{ct}}$, cross-cell-type $\rho_{\text{cross}}$, and background $\rho_{\text{bg}}$; and crosstalk probability $p_{\text{cross}}$.
2:Output: Sparse expression matrix $\mathbf{X}\in\mathbb{R}^{E\times V}$ with ground-truth labels.
3:Initialize $\mathbf{X}\leftarrow\mathbf{0}_{E\times V}$.
4:Sample cell type sizes: $g_{k}\sim\text{Poisson}(\bar{g})$, $c_{k}\sim\text{Poisson}(\bar{c})$ for $k=1,\ldots,K$.
5:Phase 1: Within-Cell-Type Signal
6:for each cell type $k=1,\ldots,K$ do
7:  For each gene $e\in E_{k}$ and cell $v\in V_{k}$: with probability $\rho_{\text{ct}}$, set $\mathbf{X}[e,v]\leftarrow\text{Poisson}(\lambda_{\text{ct}})$.
8:end for
9:Phase 2: Background Noise
10:Randomly fill $\rho_{\text{bg}}$ fraction of remaining zero entries with $\text{Poisson}(\lambda_{\text{bg}})$ values.
11:Phase 3: Cross-Cell-Type Crosstalk
12:for each cell type pair $(k,l)$ where $k\neq l$ do
13:  With probability $p_{\text{cross}}$: randomly fill $\rho_{\text{cross}}$ fraction of $\mathbf{X}[E_{k},V_{l}]$ with $\text{Poisson}(\lambda_{\text{cross}})$ values.
14:end for
15:return $\mathbf{X}$, ground-truth labels.